# A Theory of Learning with Autoregressive Chain of Thought

### v1

@@ 6 A Universal Autoregressive Base Class and Natural Emergence of Attention
In this section, we answer the optimistic question asked in the previous section, providing an explicit generator class satisfying all three desiderata. To do so, we first have to decide on a model of computation and formalize the notions of program length and runtime. We will work with a Turing machine as our model of computation, which we formalize in Section 6.1. In Section 6.2, we present our universal base class ${\mathcal{F}}_{\scriptscriptstyle{\mathsf{TM},\mathtt{S}}}$ and prove that it satisfies our desiderata from the end of the previous section (with simple input and output transformations). Finally, in Section 6.3, we discuss how attention arises naturally for simulating this class.
@@ 6.1 Turing Machine Computation
[…]
@@
The number of states of the Turing machine corresponds to program length—roughly speaking, we can think of each state corresponding to a statement in the program flow chart. More formally, the description length of a Turing machine $\langle\mathtt{S},\mathtt{T},\tau\rangle$ is $O(\mathtt{S}\log\mathtt{S})$, and by the Complexity-Theoretic Church-Turing (aka Invariance) Thesis, any reasonable model of computation can be simulated by a Turing machine up to polynomial blowup in runtime and additive overhead in description length. In particular, ${\sf PROG}(\mathtt{S},\mathtt{T})$ in the RAM model of computation satisfies ${\sf PROG}(\mathtt{S},\mathtt{T})\subseteq{\sf TM}(\tilde{O}(\mathtt{S}),{\mathcal{P}oly}(\mathtt{T}))$. Thus it suffices to address the universality question for the class ${\sf TM}(\mathtt{S},\mathtt{T})$.
@@ 6.2 Universal Autoregressive Base Class
where each $f_{\tau}$ maps a token sequence ${\bm{z}}=({\bm{z}}[\,i\,].{\mathsf{state}},{\bm{z}}[\,i\,].{\mathsf{symb}},{\bm{z}}[\,i\,].{\mathsf{move}})_{i\in|{\bm{z}}|}$ encoding the input and the history of computation up to time $t=\left\lvert{{\bm{z}}}\right\rvert-(\textnormal{input length})-1$, to the next step $z=(s_{t},a_{t},b_{t})=f_{\tau}({\bm{z}})$ the Turing machine specified by $\tau$ would make. This is specified explicitly in Algorithm 1.
@@
4  ${\mathsf{pos}}[\,i\,]=\sum_{j<i}{\bm{z}}[\,j\,].{\mathsf{move}}$triangleright Head position before a move, where ${\bm{z}}[\,i\,].{\mathsf{symb}}$ got written.
@@
We now argue that this universal autoregressive base class satisfies all three desiderata from Section 5. The definition of $f_{\tau}$ as simulating computation by a Turing machine with transition table $\tau$ directly ensures that, for any transition rule $\tau$ and input ${\bm{\omega}}$
@@
and thus ${\sf TM}(\mathtt{S},\mathtt{T})\subseteq{\mathsf{Post}}\circ{\mathcal{F}}_{\scriptscriptstyle{\mathsf{TM},\mathtt{S}}}^{{\sf e2e}\textnormal{-}{\mathtt{T}}}\circ{\mathsf{Pre}}$. Also, as there are only $(6\mathtt{S})^{3\mathtt{S}}$ possible transition tables, and so $\log\left\lvert{{\mathcal{F}}_{\scriptscriptstyle{\mathsf{TM},\mathtt{S}}}}\right\rvert=O(\mathtt{S}\log\mathtt{S})$. Applying the cardinality-based bound from Theorem 3.1, we have
@@
Given examples $({\bm{u}}_{i},v_{i})_{i=1,\ldots,\tilde{m}}$ where ${\bm{u}}_{i}\in\Sigma^{*}$ and $v_{i}\in\Sigma$ : $\text{Return $f_{\hat{\tau}}$, for some $\hat{\tau}$ with $\mathtt{S}$ states such that }\,f_{\hat{\tau}}({\bm{u}}_{i})=v_{i},\,\;\forall{i\in[\tilde{m}]}\,.$ ($\textsc{Cons}_{{\mathcal{F}}_{\scriptscriptstyle{\mathsf{TM},\mathtt{S}}}}$) Implementation: 1For each $i\in[\tilde{m}]\,:$ fill $\hat{\tau}({\mathsf{state}}_{i},{\mathsf{read}}_{i})=v_{i}$ where $({\mathsf{state}}_{i},{\mathsf{read}}_{i})=\textnormal{read-tape}({\bm{u}}_{i})\,.$ 2If there are conflicts, then there is no consistent $f_{\hat{\tau}}$. 3return $f_{\hat{\tau}}$ after completing the other entries of $\hat{\tau}(\cdot,\cdot)$ arbitrarily.
@@
For every $\mathtt{S}\in{\mathbb{N}}_{+}$ and $\mathtt{T}\geq\mathtt{S}$, we have that ${\sf TM}(\mathtt{S},\mathtt{T})$ is contained in ${\mathcal{F}}_{\scriptscriptstyle{\mathsf{TM},\mathtt{S}}}^{{\sf e2e}\textnormal{-}{\mathtt{T}}}$ (up to pre and post processing steps), and ${\mathcal{F}}_{\scriptscriptstyle{\mathsf{TM},\mathtt{S}}}^{{\sf e2e}\textnormal{-}{\mathtt{T}}}$ is ${\sf CoT}$-learnable by $\textsc{Cons}_{{\sf CoT}}$ with runtime ${\mathcal{P}oly}(n,\mathtt{T},\varepsilon^{\scriptscriptstyle{-1}},\log\delta^{\scriptscriptstyle{-1}})$ and sample complexity $m^{{\sf CoT}\textnormal{-}{\mathtt{T}}}=O\left(\varepsilon^{\scriptscriptstyle{-1}}\left(\mathtt{S}\log\mathtt{S}+\log\delta^{\scriptscriptstyle{-1}}\right)\right)\,$.
This is a corollary of the observations in the preceding paragraphs and Theorem 3.6.
@@ 6.3 Emergence of Attention
We now take a closer look at the generator class ${\mathcal{F}}_{\scriptscriptstyle{\mathsf{TM},\mathtt{S}}}$, and its generators $f_{\tau}\in{\mathcal{F}}_{\scriptscriptstyle{\mathsf{TM},\mathtt{S}}}$ specified in Algorithm 1, and how these generators can be implemented using “generic” function classes. We intentionally wrote $f_{\tau}$ as operating on arrays of length $N$. The two operations involving array elements across multiple locations are the position calculation in 4 of read-tape, which involves summing moves over multiple array locations, and the position lookup in 6 on read-tape. The lookup operation is explicitly an attention-type operation: we look for a location where the write position (the “key”) matches the read position (the “query”) and copy its tape symbol (the “value”). Meanwhile, the summation operation in 4 can be implemented as averaging many array locations, as in uniform attention.
To be more concrete, we will show how 4 and 6 can be naturally implemented by Average Hard Attention (e.g. Merrill et al., 2022, Barceló et al., 2023, Strobl et al., 2024):
@@
[…]
@@
That is, the output $\mathsf{AHA}({\mathsf{q}},{\mathsf{k}},{\mathsf{v}})[\,i\,]\in\mathbb{R}^{p}$ at position $i$ is an average of the “values” ${\mathsf{v}}[\,\ell\,]$ at positions $\ell\in\arg\max_{j\leq i}\left\langle{\mathsf{q}}[\,i\,],{\mathsf{k}}[\,j\,]\right\rangle$ whose “keys” ${\mathsf{k}}[\,j\,]$ have the maximal inner product with the “query” ${\mathsf{q}}[\,i\,]$ at the output position $i$ of interest.
Implementing 4:
@@
In the preceding paragraphs, we discussed how to implement the “non-local” operations on 4 and 6 using attention. All other operations in the implementation of $f_{\tau}$ are “local” in the sense that they operate independently at each location of the arrays [^11] This includes the operations inside read-tape which operate explicitly at array locations, and the $\tau$-lookup at the end, which can be thought of as operating at location $N$, or in fact at each location independently and in parallel outputting the computation-step-token. This directly suggest an decoder-only-transformer architecture (Vaswani, 2017) where attention operations are interleaved with multi-layer perceptrons (MLPs) operating independently and in parallel at each location, as a generic way of implementing the local operations. Most notably, the $\tau$-lookup operation is directly implementable by an MLP with $\tau$ encoded in its weights. Using Hard Average Attention as in Definition 5 results in a discontinuous and non-differentiable model, but a natural alternative is to relax the hard max in the score definition (14) to a softmax, resulting in the familiar softmax attention widely used in practice. Implementing $f_{\tau}$ with a practical transformer architecture further requires care to numerical precision, and to the specifics of implementing or approximating the local operations with MLPs and possible layer-normalizations, especially the division and multiplication operations involved in calculating the step-counters and positions. In fact, the Turing Machine expressivity results of Pérez et al. (2021), Wei et al. (2022), Merrill and Sabharwal (2023) essentially do this, following a construction similar to ours, and showing how to implement it with specific transformer architectures.
The difference here is mostly that of perspective: while recent work on Turing universality of transformers started with the transformer as suggested by Vaswani (2017) and showed how to coerce it to simulate a Turing machine, we take the reverse perspective. Our starting point is the desiderata of Section 5, which directly motivates the generator class ${\mathcal{F}}_{\scriptscriptstyle{\mathsf{TM},\mathtt{S}}}$, and we then argue how implementing $f_{\tau}\in{\mathcal{F}}_{\scriptscriptstyle{\mathsf{TM},\mathtt{S}}}$ naturally involves attention and yields a transformer-type architecture.
@@
While the tape lookup operation in 6 very naturally motivates attention, and also seems unavoidable and fundamental to universal computation, here we discuss how the use of uniform attention to calculate the current Turing Machine head position in 4 can be avoided. To do so, instead of writing out the relative move $b_{t}$ of each step of the Turing computation (as in Definition 3), we can write out the absolute tape position $p_{t}$. We cannot do so with a single token, since $p_{t}$ could potentially be as large as the runtime $\mathtt{T}$, necessitating an alphabet of size $\Sigma=\Omega(\mathtt{T})$, something we would like to avoid. But we could easily write down (i.e. generate) $p_{t}$ over multiple tokens, even over a binary alphabet $\Sigma=\{0,1\}$, using some prefix unambiguous encoding. The generator $f_{\tau}$ would then need to collect the encoding of $p_{t-1}$ (as well as the previous state and symbol written), do the arithmetic to calculate the next absolute position $p_{t}=p_{t-1}+b_{t}$ (as well other operations in Algorithm 1), figure out where in the middle of generating the description of the next compute state we are at, and generate the next bit in the description. Implementing this with a transformer-like architecture involves multiple attention layers, mostly with short-range attention to the last $O(\log\mathtt{T})$ tokens as well as a single long-term sparse attention to implement the lookup on 6, and a transformer of size (and depth) that scales with $\log\mathtt{T}$. But it avoids the need for uniform attention. A similar approach can also simulate RAM-computation, where the non-local operations are reads from memory addresses describable with $\log\mathtt{T}$ bits.
@@ 7 Discussion and Future Directions
In this paper, we presented a framework of time-invariant autoregressive Chain-of-Thought learning. We hope we convinced the reader that this is a rich and interesting framework to study. Using this framework, we discussed how:
@@
We were directly inspired by the work of Malach (2023), and address what we view as its major deficiency, namely time dependence. Our treatment is also much more comprehensive, establishing generic sample complexity results and proving hardness of ${\sf e2e}$ learning.
@@
Autoregressive generation can be viewed as a corner case of behavioral cloning (Pomerleau, 1988, Bain and Sammut, 1995), and indeed Behavioral Cloning has been suggested as an approach for understanding autoregressive generation (Chang et al., 2023, Block et al., 2023). To view autoregressive generation as a Markov Decision Process (MDP): the “state” at step $t$ is the entire sequence generated so far, i.e. $f^{{\sf CoT}\textnormal{-}{t}}({\bm{x}})$; actions correspond to tokens $a\in\Sigma$; transitions are deterministic mapping a state ${\bm{z}}$ and an action $a\in\Sigma$ to the next state $\textrm{append}({\bm{z}},a)$; and the initial state distribution corresponds to our input distribution ${\mathcal{D}}$. Our “generators” $f$ become deterministic “policies” outputting the action to take (i.e. token to append) at each state (i.e. string, or “context”). Importantly, the reward is only on the final state ${\bm{z}}=f^{{\sf CoT}\textnormal{-}{t}}({\bm{x}})$ and depends only on the correspondence of the last token in ${\bm{z}}$ to the prefix containing the initial state. Behavioral Cloning considers learning a policy maximizing the reward based only on observations from an expert, in our case the ground truth $f_{*}$, and without observing the rewards. As with much of the study of Behavioral Cloning, including the relevant recent work of Foster et al. (2024), we assume “realizability”, in that the demonstrating expert $f_{*}$ is in our policy class (i.e. base class ${\mathcal{F}}$ of generators). Behavioral Cloning always assumes the expert demonstrations are fully visible, and so corresponds to our notion of Chain-of-Thought learnability. Indeed, our learning guarantee for Chain-of-Thought learning based on the cardinality $\left\lvert{{\mathcal{F}}}\right\rvert$ (the implication of Theorem 3.1 to Chain-of-Thought learnability, though not the End-to-End learnability) is a special edge-case the learning guarantees of Foster et al. (2024). Specializing to autoregressive generation, we go well beyond cardinality-based guarantees, discuss VC-based guarantees, gaps between End-to-End and Chain-of-Thought learnability, computational hardness.
We also note that Foster et al. (2024) made a similar distinction of time invariance (parameter sharing) versus time-dependent policy classes (independent policies at each time step) in discussing Behavioral Cloning, and how this is essential for avoiding extra dependencies on the “horizon” (equivalent to our generation length).
@@
In Behavioral Cloning, the lack of reward observability makes it necessary to consider randomized policies, unlike supervised learning where deterministic policies are always optimal. Much of the emphasis in the work on Behavioral Cloning, including most of the effort and innovation of Foster et al. (2024), is in handling this stochasticity in both the policy class and expert demonstrator. But in our setting, when mapped to Behavioral Cloning, observing the expert demonstration already tells us what the reward is, and so this point is somewhat moot.
Other Issues and Open Questions.
@@
Our work leaves several technical questions open. What is the true sample complexity of ${\sf e2e}$ and ${\sf CoT}$ learning of iterated linear thresholds, i.e. lower bounds for Corollary 4.2? Can the circuit embedding construction of Lemma 4.5 be improved? Can the dependencies on $\log T$ in Table 1 be resolved, and more significantly the dependence on $\log\left\lvert{\Sigma}\right\rvert$ mentioned in Remark 3.1? Can ${\sf e2e}$ learnability be ensured using a notion more relaxed than the Littlestone dimension?
@@
Additionally, it is interesting to study other base classes under our framework, including sparse linear thresholds, gated generators, simple attention mechanisms and transformers. What are the sample and computational complexities of ${\sf e2e}$ and ${\sf CoT}$ learning? What is the expressiveness of ${\mathcal{F}}^{{\sf e2e}\textnormal{-}{T}}$? Another major challenge is generalizing our framework to the agnostic setting. In particular, what is the correct way to extend ${\sf CoT}$ learning to an agnostic setting?
@@
Our framework also allows us to rigorously study the computational complexity of settings on the spectrum between ${\sf e2e}$ and ${\sf CoT}$ learning. What can be done tractably if we have a small amount of full CoT data, and a large amount of end-to-end input-output pairs? What is the minimum amount of CoT data required for tractable learning if we have access to an unlimited amount of end-to-end data? How helpful is CoT data from a mixture of different generators $f_{i}$ all computing the same end-to-end function $f_{i}^{{\sf e2e}\textnormal{-}{T}}=f_{j}^{{\sf e2e}\textnormal{-}{T}}$ but $f_{i}^{{\sf CoT}\textnormal{-}{T}}\neq f_{j}^{{\sf CoT}\textnormal{-}{T}}$

### v2

@@ 6 A Universal Autoregressive Base Class and Natural Emergence of Attention
In this section, we answer the optimistic question asked in the previous section, providing an explicit generator class satisfying all three desiderata. To do so, we first have to decide on a model of computation and formalize the notions of program length and runtime. We will work with a Turing machine as our model of computation, which we formalize in Section˜6.1. In Section˜6.2, we present our universal base class ${\mathcal{F}}_{\scriptscriptstyle{\mathsf{TM},\mathtt{S}}}$ and prove that it satisfies our desiderata from the end of the previous section (with simple input and output transformations). Finally, in Section˜6.3, we discuss how attention arises naturally for simulating this class.
@@ 6.1 Turing Machine Computation
At the beginning of the computation, the tape is initialized with an input ${\bm{\omega}}\in\{0,1\}^{*}$ on the cells indexed from $1$ through $|{\bm{\omega}}|$ and the other cells contain $\mathbin{\square}\;$ . The head of the machine is at the position $p_{0}=|{\bm{\omega}}|+1$ and its state is initialized to $s_{0}=1\in[\mathtt{S}]$. During each time step $1\leq t\leq\mathtt{T}$, the machine reads the symbol $r_{t}=\text{Tape}[p_{t-1}]$ at the current head position $p_{t-1}$, and according to the transition rule $(s_{t},a_{t},b_{t})=\tau(s_{t-1},r_{t})$ updates its internal state to $s_{t}$, writes $\text{Tape}[p_{t-1}]\leftarrow a_{t}$ on the tape, and updates the head location to $p_{t}=p_{t-1}+b_{t}$ . Finally, after the completion of $\mathtt{T}$ many time steps, the machine stops111111 In a standard textbook definition of a turning machine (e.g. Sipser, 1996), there is a halting state, which we do not require. This is because we have a bounded runtime $\mathtt{T}$. W.l.o.g. one can always consider $\tau$ in a way that the tape does not change once the machine reaches either of the halting states, and the computation has effectively stopped. and outputs the symbol $a_{\mathtt{T}}$ (the symbol it wrote in the final step before moving the head). We denote the function computed by the machine as $g_{\langle\mathtt{S},\mathtt{T},\tau\rangle}({\bm{\omega}})=a_{\mathtt{T}}$.
@@
The number of states of the Turing machine corresponds to program length—roughly speaking, we can think of each state corresponding to a statement in the program flow chart. More formally, the description length of a Turing machine $\langle\mathtt{S},\mathtt{T},\tau\rangle$ is $O(\mathtt{S}\log\mathtt{S})$, and by the Complexity-Theoretic Church-Turing (aka Invariance) Thesis, any reasonable model of computation can be simulated by a Turing machine up to polynomial blowup in runtime and additive overhead in description length. In particular, ${\sf PROG}(\mathtt{S},\mathtt{T})$ in the RAM model of computation satisfies ${\sf PROG}(\mathtt{S},\mathtt{T})\subseteq{\sf TM}(\tilde{O}(\mathtt{S}),{\mathcal{P}\hskip-1.0ptoly}(\mathtt{T}))$. Thus it suffices to address the universality question for the class ${\sf TM}(\mathtt{S},\mathtt{T})$.
@@ 6.2 Universal Autoregressive Base Class
where each $f_{\tau}$ maps a token sequence ${\bm{z}}=({\bm{z}}[\,i\,].{\mathsf{state}},{\bm{z}}[\,i\,].{\mathsf{symb}},{\bm{z}}[\,i\,].{\mathsf{move}})_{i\in|{\bm{z}}|}$ encoding the input and the history of computation up to time $t=\left\lvert{{\bm{z}}}\right\rvert-(\textnormal{input length})-1$, to the next step $z=(s_{t},a_{t},b_{t})=f_{\tau}({\bm{z}})$ the Turing machine specified by $\tau$ would make. This is specified explicitly in Algorithm˜1.
@@
4  ${\mathsf{pos}}[\,i\,]=\sum_{j<i}{\bm{z}}[\,j\,].{\mathsf{move}}$ $\triangleright$ Head position before a move, where ${\bm{z}}[\,i\,].{\mathsf{symb}}$ got written.
@@
We now argue that this universal autoregressive base class satisfies all three desiderata from Section˜5. The definition of $f_{\tau}$ as simulating computation by a Turing machine with transition table $\tau$ directly ensures that, for any transition rule $\tau$ and input $\bm{\omega}$}
@@
and thus ${\sf TM}(\mathtt{S},\mathtt{T})\subseteq{\mathsf{Post}}\circ{\mathcal{F}}_{\scriptscriptstyle{\mathsf{TM},\mathtt{S}}}^{{\sf e2e}\textnormal{-}{\mathtt{T}}}\circ{\mathsf{Pre}}$. Also, as there are only $(6\mathtt{S})^{3\mathtt{S}}$ possible transition tables, and so $\log\left\lvert{{\mathcal{F}}_{\scriptscriptstyle{\mathsf{TM},\mathtt{S}}}}\right\rvert=O(\mathtt{S}\log\mathtt{S})$. Applying the cardinality-based bound from Theorem˜3.1, we have
@@
[…]
@@
For every $\mathtt{S}\in{\mathbb{N}}_{+}$ and $\mathtt{T}\geq\mathtt{S}$, we have that ${\sf TM}(\mathtt{S},\mathtt{T})$ is contained in ${\mathcal{F}}_{\scriptscriptstyle{\mathsf{TM},\mathtt{S}}}^{{\sf e2e}\textnormal{-}{\mathtt{T}}}$ (up to pre and post processing steps), and ${\mathcal{F}}_{\scriptscriptstyle{\mathsf{TM},\mathtt{S}}}^{{\sf e2e}\textnormal{-}{\mathtt{T}}}$ is $\sf CoT$}-learnable by $\textsc{Cons}_{{\sf CoT}}$ with runtime ${\mathcal{P}\hskip-1.0ptoly}(n,\mathtt{T},\varepsilon^{\scriptscriptstyle{-1}},\log\delta^{\scriptscriptstyle{-1}})$ and sample complexity $m^{{\sf CoT}\textnormal{-}{\mathtt{T}}}=O\left(\varepsilon^{\scriptscriptstyle{-1}}\left(\mathtt{S}\log\mathtt{S}+\log\delta^{\scriptscriptstyle{-1}}\right)\right)\,$.
This is a corollary of the observations in the preceding paragraphs and Theorem˜3.6.
@@ 6.3 Emergence of Attention
We now take a closer look at the generator class ${\mathcal{F}}_{\scriptscriptstyle{\mathsf{TM},\mathtt{S}}}$, and its generators $f_{\tau}\in{\mathcal{F}}_{\scriptscriptstyle{\mathsf{TM},\mathtt{S}}}$ specified in Algorithm˜1, and how these generators can be implemented using “generic” function classes. We intentionally wrote $f_{\tau}$ as operating on arrays of length $N$. The two operations involving array elements across multiple locations are the position calculation in ˜4 of read-tape, which involves summing moves over multiple array locations, and the position lookup in ˜6 on read-tape. The lookup operation is explicitly an attention-type operation: we look for a location where the write position (the “key”) matches the read position (the “query”) and copy its tape symbol (the “value”). Meanwhile, the summation operation in ˜4 can be implemented as averaging many array locations, as in uniform attention.
To be more concrete, we will show how ˜4 and 6 can be naturally implemented by Average Hard Attention (e.g. Merrill et al., 2022, Barceló et al., 2023, Strobl et al., 2024):
@@
For any positive integers $N$, $p$, $l$ and any sequences of vectors ${\mathsf{q}}[\,1\,],\ldots,{\mathsf{q}}[N]\in\mathbb{R}^{l}$ (queries), ${\mathsf{k}}[\,1\,],\ldots,{\mathsf{k}}[N]\in\mathbb{R}^{l}$ (keys), and ${\mathsf{v}}[\,1\,],\ldots,{\mathsf{v}}[N]\in\mathbb{R}^{p}$ (values), and a score function $\mathsf{score}:\mathbb{R}^{i}\to\mathbb{R}^{i}$ for each $i\in[N]$, the causal attention function $\mathsf{Att}:(\mathbb{R}^{l})^{N}\times(\mathbb{R}^{l})^{N}\times(\mathbb{R}^{p})^{N}\mapsto(\mathbb{R}^{p})^{N}$ is defined as:
@@
That is, the output $\mathsf{AHA}({\mathsf{q}},{\mathsf{k}},{\mathsf{v}})[\,i\,]\in\mathbb{R}^{p}$ at position $i$ is an average of the “values” $\mathsf{v}$$\ell\,$] at positions $\ell\in\arg\max_{j\leq i}\left\langle{\mathsf{q}}[\,i\,],{\mathsf{k}}[\,j\,]\right\rangle$ whose “keys” ${\mathsf{k}}[\,j\,]$ have the maximal inner product with the “query” ${\mathsf{q}}[\,i\,]$ at the output position $i$ of interest.
Implementing ˜4:
@@
In the preceding paragraphs, we discussed how to implement the “non-local” operations on ˜4 and 6 using attention. All other operations in the implementation of $f_{\tau}$ are “local” in the sense that they operate independently at each location of the arrays [^12] This includes the operations inside read-tape which operate explicitly at array locations, and the $\tau$-lookup at the end, which can be thought of as operating at location $N$, or in fact at each location independently and in parallel outputting the computation-step-token. This directly suggests a decoder-only-transformer architecture (Vaswani, 2017) where attention operations are interleaved with multi-layer perceptrons (MLPs) operating independently and in parallel at each location, as a generic way of implementing the local operations. Most notably, the $\tau$-lookup operation is directly implementable by an MLP with $\tau$ encoded in its weights. Using Hard Average Attention as in Definition˜5 results in a discontinuous and non-differentiable model, but a natural alternative is to relax the hard max in the score definition (14) to a softmax, resulting in the familiar softmax attention widely used in practice. Implementing $f_{\tau}$ with a practical transformer architecture further requires care to numerical precision, and to the specifics of implementing or approximating the local operations with MLPs and possible layer-normalizations, especially the division and multiplication operations involved in calculating the step-counters and positions. In fact, the Turing Machine expressivity results of Pérez et al. (2021), Wei et al. (2022), Merrill and Sabharwal (2023) essentially do this, following a construction similar to ours, and showing how to implement it with specific transformer architectures.
The difference here is mostly that of perspective: while recent work on Turing universality of transformers started with the transformer as suggested by Vaswani (2017) and showed how to coerce it to simulate a Turing machine, we take the reverse perspective. Our starting point is the desiderata of Section˜5, which directly motivates the generator class ${\mathcal{F}}_{\scriptscriptstyle{\mathsf{TM},\mathtt{S}}}$, and we then argue how implementing $f_{\tau}\in{\mathcal{F}}_{\scriptscriptstyle{\mathsf{TM},\mathtt{S}}}$ naturally involves attention and yields a transformer-type architecture.
@@
While the tape lookup operation in ˜6 very naturally motivates attention, and also seems unavoidable and fundamental to universal computation, here we discuss how the use of uniform attention to calculate the current Turing Machine head position in ˜4 can be avoided. To do so, instead of writing out the relative move $b_{t}$ of each step of the Turing computation (as in Definition˜3), we can write out the absolute tape position $p_{t}$. We cannot do so with a single token, since $p_{t}$ could potentially be as large as the runtime $\mathtt{T}$, necessitating an alphabet of size $\Sigma=\Omega(\mathtt{T})$, something we would like to avoid. But we could easily write down (i.e. generate) $p_{t}$ over multiple tokens, even over a binary alphabet $\Sigma=\{0,1\}$, using some prefix unambiguous encoding. The generator $f_{\tau}$ would then need to collect the encoding of $p_{t-1}$ (as well as the previous state and symbol written), do the arithmetic to calculate the next absolute position $p_{t}=p_{t-1}+b_{t}$ (as well other operations in Algorithm˜1), figure out where in the middle of generating the description of the next compute state we are at, and generate the next bit in the description. Implementing this with a transformer-like architecture involves multiple attention layers, mostly with short-range attention to the last $O(\log\mathtt{T})$ tokens as well as a single long-term sparse attention to implement the lookup on ˜6, and a transformer of size (and depth) that scales with $\log\mathtt{T}$. But it avoids the need for uniform attention. A similar approach can also simulate RAM-computation, where the non-local operations are reads from memory addresses describable with $\log\mathtt{T}$ bits.
@@ 7 Discussion and Future Directions
We investigate training based solely on information about the final answer, as well as training with supervision about the intermediate reasoning steps used to reach that answer. Training based on explicit Chain-of-Thought examples is studied by Nye et al. (2021), Zelikman et al. (2022), Chung et al. (2024), Lewkowycz et al. (2022), and is widely used in contemporary training of frontier LLMs, particularly in the fine-tuning step (OpenAI, 2023, DeepMind, 2024, Touvron et al., 2023).
More recently, there have been attempts to train models using only feedback on the final answer, attempting to learn the reasoning process using reinforcement learning techniques (Guo et al., 2025, Ouyang et al., 2022). Our results help highlight and formalize the intractability of such approaches in general, without additional assumptions. So, why have approaches like DeepSeek and others seen empirical success? One possible explanation is that these Reinforce-type methods assume access to a sufficiently good (possibly randomized) next-token predictor $f\in\mathcal{F}$ that already produces correct (or at least useful) CoTs with non-negligible probability with some level of coverage over valid reasoning paths. Such a predictor is typically learned during pretraining on large-scale text data, which, while not always matching the target task’s CoTs exactly, are often similar enough to enable transfer. This type of training is not exactly CoT supervision, but is arguably closer to it than pure end-to-end supervision. Understanding how transfer from such related-but-different next-token prediction tasks enables tractable learning—using only final-answer supervision, and little or no direct CoT data for the task of interest—is an extremely interesting direction for future work.
In summary, we presented a framework of time-invariant autoregressive Chain-of-Thought learning. Using our framework, we discussed how:
@@
We were directly inspired by the work of Malach (2023), and address what we view as its major deficiency, namely time dependence.
@@
Recent work has connected autoregressive generation to Behavior Cloning (BC) (Pomerleau, 1988, Bain and Sammut, 1995, Ross and Bagnell, 2010), and indeed BC has been suggested as an approach for understanding modern language models (Chang et al., 2023, Block et al., 2023). Those works view autoregressive generation as a Markov Decision Process (MDP) with a “state” at step $t$ being the entire sequence generated thus far, i.e. $f^{{\sf CoT}\textnormal{-}{t}}({\bm{x}})$, actions corresponding to tokens $a\in\Sigma$, and transitions being deterministic mapping a state $\bm{z}$} and an action $a\in\Sigma$ to the next state $\textrm{append}({\bm{z}},a)$. In that framework, our “generators” $f$ become deterministic “policies” outputting the action to take (i.e. token to append) at each state (i.e. string, or “context”). Importantly, the reward is only on the final state ${\bm{z}}=f^{{\sf CoT}\textnormal{-}{t}}({\bm{x}})$ and depends only on the correspondence of the last token in $\bm{z}$} to the prefix containing the initial state. BC considers learning a policy maximizing some reward based only on observations from an expert, in our case the ground truth $f_{*}$, and without observing the ground-truth rewards. Most relevant to our paper is the recent work of Foster et al. (2024), which provides general, horizon-free (i.e., independent of $T$) guarantees on the performance of BC. As in our work, Foster et al. (2024) considers the realizable setting, where the expert $f_{*}$ is in a given policy class, and distinguishes between the time-invariant (called parameter-sharing in that work) and time-varying cases, noting that the former is essential for the aforementioned horizon-free guarantees. Foster et al. (2024) considers only the CoT framework, where ‘actions’ at each time step are available to the learner, and indeed, the implication of Theorem˜3.1 for CoT learning is a special case of their result. That work is focused on the statistical task itself and is not interested in either the end-to-end setting or the computational benefits and drawbacks thereof.
@@
Another innovation in Foster et al. (2024) is the consideration of stochasticity in both the environment and the experts. The lack of reward observability (combined with a stochastic expert) necessitates the use of randomized policies—unlike in supervised learning, where deterministic policies are always optimal. In our setting, when mapped to Behavioral Cloning, observing the expert demonstration already reveals the reward, so this consideration is somewhat moot.
Further Open Questions.
@@
Our work leaves several technical questions open such as: What is the true sample complexity of $\sf e2e$} and $\sf CoT$} learning of iterated linear thresholds, i.e. lower bounds for Corollary˜4.2? Can the circuit embedding construction of Lemma˜4.5 be improved? Can the dependencies on $\log T$ in Table˜1 be resolved, and more significantly the dependence on $\log\left\lvert{\Sigma}\right\rvert$ mentioned in Remark˜3.1? Can $\sf e2e$} learnability be ensured using a notion more relaxed than the Littlestone dimension?
@@
Additionally, it is interesting to study other base classes under our framework, including sparse linear thresholds, gated generators, simple attention mechanisms and transformers. What are the sample and computational complexities of $\sf e2e$} and $\sf CoT$} learning? What is the expressiveness of ${\mathcal{F}}^{{\sf e2e}\textnormal{-}{T}}$? Another major challenge is generalizing our framework to the agnostic setting. In particular, what is the correct way to model misspecification for the $\sf CoT$} data?
@@
Our framework also allows us to rigorously study the computational complexity of settings on the spectrum between $\sf e2e$} and $\sf CoT$} learning. What can be done tractably if we have a small amount of full CoT data, and a large amount of end-to-end input-output pairs? What is the minimum amount of CoT data required for tractable learning if we have access to an unlimited amount of end-to-end data? How helpful is CoT data from a mixture of different generators $f_{i}$ all computing the same end-to-end function $f_{i}^{{\sf e2e}\textnormal{-}{T}}=f_{j}^{{\sf e2e}\textnormal{-}{T}}$ but $f_{i}^{{\sf CoT}\textnormal{-}{T}}\neq f_{j}^{{\sf CoT}\textnormal{-}{T}}$ ?